# The Koszul Property for Truncations of Nonstandard Graded Polynomial Rings

### v1

@@ 1. Introduction
If $S$ is a nonstandard graded polynomial ring and $S^{(e)}$ the $e$-th Veronese subring, then $S^{(e)}$ is a nonstandard Koszul algebra.
We emphasize that Theorem A alone does not lead to a proof of this conjecture. But it demonstrates that Koszulness results from the classical graded setting have analogues in the nonstandard graded settings, providing some motivation for the conjecture.
@@ Acknowledgments
We are tremendously grateful to Daniel Erman for his guidance, support, and feedback throughout this project. We would also like to thank Maya Banks, John Cobb, Jose Israel Rodriguez, and Aleksandra Sobieska for their mentorship and helpful comments. Multiple discussions with Srikanth Iyengar helped shape our understanding of Koszul modules and relevant results. This paper benefited greatly from many conversations that occurred at the Introductory Workshop in Commutative Algebra at SLMath. In particular, we are very grateful to Aldo Conca, Lauren Cranton Heller, David Eisenbud, Srikanth Iyengar, Jason McCullough, Gregory Smith, and many more for their valuable insights and suggestions. We also thank SLMath for hosting events like this that provide the opportunity for researchers to discuss their work. We acknowledge the support from the National Science Foundation Grant DMS-2200469, as well as that from the math departments at the University of Hawai’i at Manoa and the University of Wisconsin-Madison. Lastly, most computations were performed with the aid of Macaulay2 [GS].

### v2

@@ 1. Introduction
If $S$ is a nonstandard graded polynomial ring and $S^{(e)}$ the $e$-th Veronese subring, then $S^{(e)}$ is a nonstandard Koszul algebra for $e$ sufficiently large.
Remark 1.6.
We are grateful to Alexandra Seceleanu for pointing out an error in an earlier draft. We had previously excluded the hypothesis that $e$ be sufficiently large; however, examples such as $k[x,y,z]^{(15)}$ where $\deg(x)=3$, $\deg(y)=4$, and $\deg(z)=5$ prove this to be false. To motivate this additional hypothesis, note that if $R$ is any standard graded $k$-algebra, then $R^{(e)}$ is Koszul for $e$ sufficiently large.
We emphasize that Theorem A alone does not lead to a proof of this conjecture. But it demonstrates that Koszulness results from the classical graded setting have analogues in the nonstandard graded setting, providing some motivation for the conjecture.
@@ Acknowledgments
We are tremendously grateful to Daniel Erman for his guidance, support, and feedback throughout this project. We would also like to thank Maya Banks, John Cobb, Jose Israel Rodriguez, and Aleksandra Sobieska for their mentorship and helpful comments. Multiple discussions with Srikanth Iyengar helped shape our understanding of Koszul modules and relevant results, for which we are very appreciative. We are very grateful to Aldo Conca, Lauren Cranton Heller, David Eisenbud, Jason McCullough, Alexandra Seceleanu, Gregory Smith, and many more for their valuable insights and suggestions. Many of these conversations occurred at the Introductory Workshop in Commutative Algebra at SLMath, so we would also like to thank SLMath for hosting events like this that provide the opportunity for researchers to discuss their work. We acknowledge the support from the National Science Foundation Grant DMS-2200469, as well as that from the math departments at the University of Hawai‘i at Mānoa and the University of Wisconsin-Madison. Lastly, most computations were performed with the aid of Macaulay2 [GS].